# Feiyu Xiong3, Tong Xu2, Matthew B. Blaschko1

### v2

@@ Appendix B Prompt
Appendix C Generation Examples
Examples of model responses are presented in Figs. 10, 11, 13 and 12. Overall, FastMem can improve the model’s context awareness without losing general language ability. We have manually evaluated a few hundred responses and observed that although FastMem adapts the model to a small text snippet, there is no noticeable difference in the fluency and coherence of the generation results before and after applying FastMem. Additionally, we observe that the responses produced by FastMem remain as concise as the original output, while FastMem enables the model to more accurately follow the instructions and references. For knowledge-conflicting datasets such as NQ-SWAP, the constructed examples may not be proper, especially when the swapped entity is time. We observe that FastMem makes the model more aware of the context without blindly extracting the answer string.

### v3

@@ Appendix C Prompt
Appendix D Generation Examples
Examples of model responses are presented in Figs. 13, 14, 16 and 15. Overall, FastMem can improve the model’s context awareness without losing general language ability. We have manually evaluated a few hundred responses and observed that although FastMem adapts the model to a small text snippet, there is no noticeable difference in the fluency and coherence of the generation results before and after applying FastMem. Additionally, we observe that the responses produced by FastMem remain as concise as the original output, while FastMem enables the model to more accurately follow the instructions and references. For knowledge-conflicting datasets such as NQ-SWAP, the constructed examples may not be proper, especially when the swapped entity is time. We observe that FastMem makes the model more aware of the context without blindly extracting the answer string.
Figure 10: Llama 3-8B-Instruct: The number of epochs vs. performance in MemoTrap.
Figure 11: Qwen 1.5-7B-Chat: The number of epochs vs. performance in MemoTrap.